OpenStudy (anonymous):

True or false: The range of f(x)=log(x+2) is y is greater than or equal to 0

jimthompson5910 (jim_thompson5910):

This is false. If f(x) = -1, then f(x)=log(x+2) -1=log(x+2) 10^(-1) = x+2 0.1 = x+2 0.1 - 2 = x -1.9 = x x = -1.9 So because f(-1.9) = -1, we know that f(x) = -1 is indeed in the range. So that proves that the range is NOT y >= 0

OpenStudy (anonymous):

Thank you! =)

jimthompson5910 (jim_thompson5910):

you're welcome